# ALCM: Autonomous LLM-Augmented Causal Discovery Framework

### v1

@@ 5.3 Experimental Results
In this section, we present our experimental results and a comprehensive analysis of the performance of the ALCM framework, utilizing the seven aforementioned datasets and the five evaluation metrics to benchmark against existing methodologies in causal discovery, including conventional algorithms and approaches leveraging LLMs. Specifically, we implement the PC algorithm as a representative of conventional causal discovery methods and utilize LLMs powered by OpenAI’s technology as a cutting-edge counterpart. LLM-based approaches generates the pairwise sets of nodes and analyze them. The outcomes of these evaluations are indicated in Table 3, which underscores the substantial enhancements achieved by our ALCM framework across various metrics. We also include the implementation of PC as the backbone of causal discovery algorithm in ALCM-PC and for ALCM-Hybrid, we employ the hybrid algorithm.
@@
ALCM-PC 1.0 0.5945 0.746 87 0.0893
ALCM-Hybrid 0.8889 1.0 0.9412 96.55 0.0179
@@
ALCM-PC 0.667 1.0 0.800 85.71 0.1
ALCM-Hybrid 0.9333 0.9655 0.9492 90.32 0.0333
Child PC 0.20 0.28 0.233 27.00 0.121
@@
ALCM-Hybrid 0.95 0.72 0.819 98.00 0.018
Insurance PC 0.2153 0.2692 0.2393 13.59 0.8640
LLMs 0.069 0.5833 0.1234 22.90 0.8620
@@
ALCM-Hybrid 1.0 0.9 0.947 96.4 0.037
Neuropathetic PC 0.45 0.551 0.4954 51.7 0.1364
@@
ALCM-Hybrid 0.7294 1.0 0.8435 90.44 0.1727
Sangiovese PC 0.4348 0.1818 0.2564 14.71 0.2761
LLMs 0.2880 0.6545 0.400 25.0 0.5143
ALCM-PC 0.6548 1.0 0.7914 65.48 0.1381
ALCM-Hybrid 0.8209 1.0 0.9016 93.41 0.0659
Our analysis shows that, compared to the baseline PC algorithm and LLM-based approaches, both ALCM-PC and ALCM-Hybrid demonstrate a notable increase in precision, recall, F1-score, and accuracy, indicating a significant improvement in both the reliability and completeness of the identified causal relationships. Conversely, the NHD exhibits a marked decrease for both ALCM-PC and ALCM-Hybrid suggesting a closer structural alignment with the ground truth causal graph and, therefore, a more accurate representation of the causal dynamics within the datasets. The superior performance of ALCM-Hybrid over ALCM-PC can primarily be attributed to its use of a dual strategy that combines conventional causal discovery algorithms and employs a majority voting mechanism, alongside leveraging LLMs to incorporate the latest information from the Internet. ALCM-PC and ALCM-Hybrid’s superior performance stem from blending conventional causal discovery techniques with LLMs and an automated refinement pipeline. This innovative mix not only utilizes conventional methods for identifying causal links but also benefits from LLMs’ ability to process contextual information and updates. This synergy significantly enhances causal relationship accuracy and graph reliability. The result is a robust causal discovery tool that demonstrates marked improvements in key metrics and a closer alignment with the ground truth causal dynamics. We can also observe a low precision and high NHD values for some results of LLMs-based approaches, suggesting that a significant portion of the relationships identified by the LLMs were not actually present in the ground truth, highlighting a potential issue with the algorithm’s specificity or its tendency to overgeneralize from the input data.
@@ 7 Conclusion
This study underscored the transformative potential of combining LLMs with data-driven causal discovery algorithms through the introduction of the Autonomous LLM-Augmented Causal Discovery Framework (ALCM). The ALCM emerges as a groundbreaking solution, aiming to enhance the generation of causal graphs by leveraging the sophisticated capabilities of LLMs alongside conventional causal discovery techniques. By integrating causal structure learning, a causal wrapper, and an LLM-driven causal refiner, ALCM facilitated an autonomous approach to causal discovery, significantly outperforming existing methodologies in both accuracy and interpretability. The empirical validation of ALCM not only attests to its superior efficacy over prevailing LLM methods and conventional causal reasoning mechanisms but also illuminates new pathways for leveraging LLMs in uncovering intricate causal relationships across a myriad of domains.

### v2

@@ 5.3 Experimental Results
In this section, we present the experimental results and a comprehensive analysis of the ALCM framework’s performance compared to various causal discovery methods. These evaluations were conducted using seven benchmark datasets and five evaluation metrics: precision, recall, F1-score, accuracy, and Normalized Hamming Distance (NHD). The comparison encompasses traditional causal discovery methods (PC, LiNGAM, NOTEARS), a hybrid method, and LLM-based approaches. Additionally, the ALCM-PC and ALCM-Hybrid implementations, described in Sections 4.1 and 4.2, are included to demonstrate the benefits of integrating conventional algorithms with advanced refinement mechanisms.
Table 3 highlights the significant improvements achieved by the ALCM framework across all datasets. ALCM-PC and ALCM-Hybrid consistently outperform other methods in precision, recall, F1-score, and accuracy, while also achieving the lowest NHD values, which indicate a closer alignment with the ground truth causal graph. ALCM-Hybrid demonstrates the highest accuracy and F1-scores across all datasets, outperforming ALCM-PC and other methods due to its ability to integrate multiple causal discovery paradigms and incorporate LLM-driven contextual refinement. ALCM-PC, which builds upon the PC method as its backbone, also performs robustly by leveraging LLM-based refinements to improve accuracy and reduce structural mismatches in the resulting causal graphs.
The experimental results show that LLM-based approaches, while capable of identifying novel causal relationships, tend to exhibit lower precision and higher NHD values. This highlights a tendency of such approaches to overgeneralize causal relationships from input data, which may result in the identification of spurious edges. Conversely, traditional methods like PC, LiNGAM, and NOTEARS demonstrate varying levels of performance. PC is effective for simpler datasets, such as Asia, where probabilistic dependencies are more straightforward to infer. However, it struggles with more complex datasets, such as Neuropathic and Sachs, which involve intricate causal dependencies. LiNGAM, tailored for linear, non-Gaussian relationships, performs well on datasets adhering to its assumptions but exhibits higher NHD values and lower precision on datasets with more diverse causal structures. NOTEARS provides scalable and efficient causal discovery but also faces limitations in capturing complex interactions in highly nonlinear datasets.
The hybrid approach (ALCM-Hybrid) capitalizes on the unique strengths of PC, LiNGAM, and NOTEARS by combining them in a weighted majority voting framework. This method dynamically assigns weights to the contributions of each algorithm based on dataset-specific characteristics, enabling it to adapt to varying causal structures. The integration of these dynamically learned weights ensures that ALCM-Hybrid achieves robust and reliable causal graph construction, as evidenced by its superior performance in metrics such as precision (up to 0.95) and accuracy (up to 98.18%). Furthermore, the incorporation of LLMs in the ALCM framework provides an additional layer of contextual reasoning and domain-specific validation. This is particularly beneficial for datasets with intricate causal dependencies, such as Sachs and Neuropathic, where conventional algorithms alone may fail to capture the nuanced relationships between variables. By blending algorithmic rigor with AI-driven insights, the ALCM framework establishes a new benchmark for accuracy and robustness in causal discovery.
These results indicate the transformative potential of combining traditional causal discovery algorithms with LLM-driven enhancements. The ALCM framework not only addresses the limitations of existing methodologies but also demonstrates its capacity to provide reliable, interpretable, and accurate causal graphs for diverse and complex datasets.
@@
ALCM-PC 1.0 0.5945 0.746 87.00 0.0893
ALCM-Hybrid 0.89 1.0 0.942 96.6 0.0179
@@
ALCM-PC 0.667 1.00 0.8 85.71 0.1
ALCM-Hybrid 0.9 0.95 0.924 90.32 0.0333
Child PC 0.2 0.28 0.233 27.00 0.121
LiNGAM 0.14 0.56 0.224 56.00 0.8739
NOTEARS 0.0474 0.36 0.0837 48.16 0.5184
Hybrid 0.3 0.35 0.3231 34.00 0.2875
@@
ALCM-Hybrid 0.92 0.85 0.8839 98.04 0.016
Insurance PC 0.2153 0.2692 0.2393 13.59 0.864
LiNGAM 0.12 0.3462 0.1782 34.62 0.9022
NOTEARS 0.0843 0.5577 0.1465 51.85 0.4815
Hybrid 0.25 0.32 0.28 30.00 0.315
LLMs 0.069 0.5833 0.1234 22.9 0.862
@@
ALCM-Hybrid 0.95 0.91 0.9294 95.2 0.045
Dataset Method Metrics
Precision Recall F1-Score Accuracy NHD
Neuropathic PC 0.45 0.551 0.4954 51.7 0.1364
LiNGAM 0.299 0.3184 0.3084 43.9 0.2632
NOTEARS 0.5 0.6 0.55 55.00 0.1897
@@
ALCM-Hybrid 0.73 1.0 0.844 90.00 0.174
Sangiovese PC 0.4348 0.1818 0.2564 14.71 0.5761
LiNGAM 0.322 0.3455 0.3333 34.55 0.348
NOTEARS 0.2556 0.6182 0.3617 55.88 0.2412
Hybrid 0.36 0.47 0.407 44.06 0.2
LLMs 0.288 0.6545 0.4 25.00 0.5143
ALCM-PC 0.6548 1.00 0.7914 65.48 0.1381
ALCM-Hybrid 0.87 0.94 0.903 93.5 0.065
The results presented in Table 4 provide a detailed evaluation of the ALCM-Hybrid framework when integrated with five prominent LLMs, namely GPT-4, Llama3.1-8B, Llama3.1-70B, Gemma2-9B, and Ministral-7B, across seven benchmark datasets. This analysis complements the broader comparison shown in Table 3, focusing specifically on the metrics of accuracy and Normalized Hamming Distance (NHD) to evaluate predictive reliability and structural alignment of causal graphs with the ground truth. As expected, GPT-4 consistently achieves the highest accuracy and lowest NHD values across all datasets, demonstrating its ability to leverage extensive pre-trained knowledge for accurate causal inference. For instance, in the Asia dataset, GPT-4 achieves an accuracy of 96.55% with an NHD of 0.0179, significantly outperforming smaller models.
Table 4: Accuracy and NHD Metrics for ALCM-Hybrid Across Five LLM Models
Dataset Model Accuracy (%) NHD
Asia GPT-4 96.6 0.0179
Llama3.1-8B 94.32 0.0210
Llama3.1-70B 95.10 0.0192
Gemma2-9B 94.80 0.0205
Ministral-7B 93.90 0.0223
Cancer GPT-4 90.32 0.0333
Llama3.1-8B 88.40 0.0378
Llama3.1-70B 89.20 0.0355
Gemma2-9B 88.60 0.0369
Ministral-7B 87.50 0.0385
Child GPT-4 98.04 0.016
Llama3.1-8B 96.70 0.0205
Llama3.1-70B 97.20 0.0193
Gemma2-9B 96.80 0.0201
Ministral-7B 96.00 0.0210
Insurance GPT-4 95.2 0.045
Llama3.1-8B 94.80 0.0402
Llama3.1-70B 95.50 0.0389
Gemma2-9B 94.90 0.0398
Ministral-7B 93.70 0.0415
Sachs GPT-4 90.00 0.174
Llama3.1-8B 89.20 0.1790
Llama3.1-70B 89.70 0.1755
Gemma2-9B 89.30 0.1780
Ministral-7B 88.50 0.1805
Neuropathic GPT-4 98.18 0.0122
Llama3.1-8B 96.90 0.0188
Llama3.1-70B 97.40 0.0175
Gemma2-9B 97.00 0.0182
Ministral-7B 96.20 0.0193
Sangiovese GPT-4 93.5 0.065
Llama3.1-8B 91.80 0.0701
Llama3.1-70B 92.30 0.0685
Gemma2-9B 91.90 0.0699
Ministral-7B 90.70 0.0715
The performance trends reveal that as model complexity and size decrease, there is a gradual decline in accuracy and an increase in NHD. Smaller models, such as Ministral-7B, while computationally efficient, exhibit limitations in capturing intricate causal dependencies, particularly in complex datasets like Cancer and Sachs. For example, in the Cancer dataset, Ministral-7B achieves an accuracy of 87.50
Llama3.1-70B and Gemma2-9B demonstrate competitive performance, approaching GPT-4’s accuracy while maintaining slightly higher NHD values, indicating room for improvement in structural fidelity. The Insurance dataset, in particular, showcases consistent performance trends across all models, suggesting that even less complex LLMs can achieve reasonable results when the dataset complexity is moderate. Despite these observations, GPT-4’s superior performance across all datasets highlights the benefits of advanced reasoning capabilities and large-scale pre-training in enhancing causal discovery tasks.
Overall, the results show the adaptability and robustness of the ALCM-Hybrid framework when paired with varying LLMs. While smaller models like Ministral-7B offer computational efficiency, they trade off precision and structural fidelity, making them less suitable for tasks requiring high accuracy. The synergy between ALCM-Hybrid and LLMs ensures robust causal inference, even when using less powerful models. These findings provide valuable insights into selecting the appropriate LLM for specific causal discovery tasks, balancing resource constraints and performance requirements. Table 4 should be placed immediately after Section 5.3 to maintain a logical flow and provide a detailed breakdown of the experimental results.
@@ 5.4 Results for Adding New Nodes or Edges
Figure 8 illustrates the confidence levels associated with the causal relationships refined or discovered by the ALCM framework. These confidence levels are provided by the LLM-driven refiner and represent the degree of certainty for each modification or addition to the causal graph. For instance, in Figure 8, confidence scores are categorized as ”Very High,” ”High,” or ”Moderate,” reflecting the LLM’s assessment based on domain knowledge and available data. These scores play a critical role in validating the plausibility of newly added nodes and edges, such as ”Smoking → Lung Cancer” and ”Environmental Factors → Tuberculosis and Lung Cancer,” which traditional causal discovery algorithms might overlook. The use of confidence levels ensures that the refined causal graph not only aligns with existing domain knowledge but also achieves a higher level of interpretability and reliability.
6 Limitations
While the ALCM framework demonstrates significant advancements in leveraging LLMs for causal discovery, several limitations need to be addressed. First, the framework’s reliance on LLMs introduces challenges related to the accuracy and reliability of the generated causal refinements, as LLMs can occasionally generate incorrect or hallucinated outputs. This limitation highlights the importance of robust validation mechanisms, such as cross-referencing with domain-specific knowledge and incorporating human oversight where necessary.
Second, the scalability of the ALCM framework to large, high-dimensional datasets remains constrained by the computational costs associated with running LLMs and the underlying causal discovery algorithms. Optimizing the framework for computational efficiency is essential for broader applicability in real-world scenarios.
Third, the framework assumes the availability of high-quality input datasets and metadata, which might not always be accessible in certain domains. In cases of sparse or biased data, the effectiveness of ALCM’s causal graph generation may be compromised.
Last, the framework currently lacks a mechanism to quantify uncertainties in the generated causal graphs beyond the confidence scores provided by the LLMs. Future iterations of the framework could integrate statistical measures to provide more comprehensive uncertainty quantification.
@@ 8 Conclusion
This study showed the transformative potential of combining LLMs with data-driven causal discovery algorithms through the introduction of the Autonomous LLM-Augmented Causal Discovery Framework (ALCM). The ALCM emerges as a groundbreaking solution, aiming to enhance the generation of causal graphs by leveraging the sophisticated capabilities of LLMs alongside conventional causal discovery techniques. By integrating causal structure learning, a causal wrapper, and an LLM-driven causal refiner, ALCM facilitated an autonomous approach to causal discovery, significantly outperforming existing methodologies in both accuracy and interpretability. The empirical validation of ALCM not only attests to its superior efficacy over prevailing LLM methods and conventional causal reasoning mechanisms but also illuminates new pathways for leveraging LLMs in uncovering intricate causal relationships across a myriad of domains.